# Learning to Explore and Select for Coverage-Conditioned Retrieval-Augmented Generation

### v1

@@ Abstract
Interactions with billion-scale large language models typically yield long-form responses due to their extensive parametric capacities, along with retrieval-augmented features. While detailed responses provide insightful viewpoint of a specific subject, they frequently generate redundant and less engaging content that does not meet user interests. In this work, we focus on the role of query outlining (i.e., selected sequence of queries) in scenarios that users request a specific range of information, namely coverage-conditioned (C2) scenarios. For simulating C2 scenarios, we construct QTree, 10K sets of information-seeking queries decomposed with various perspectives on certain topics. By utilizing QTree, we train QPlanner, a 7B language model generating customized query outlines that follow coverage-conditioned queries. We analyze the effectiveness of generated outlines through automatic and human evaluation, targeting on retrieval-augmented generation (RAG). Moreover, the experimental results demonstrate that QPlanner with alignment training can further provide outlines satisfying diverse user interests. Our resources are available at https://github.com/youngerous/qtree.
Learning to Explore and Select for Coverage-Conditioned
Retrieval-Augmented Generation
@@ 1 Introduction
Curiosity is attributed to perceptual knowledge and sequence of interests (BERLYNE, 1954). The emerging performance of recent large language models (LLMs) enables providing long and detailed responses by utilizing their deep parametric knowledge. Meanwhile, human-machine interaction interface (e.g., chat and acoustic interface) that has been studied for a long time (Levin et al., 2000) allows users to request customized information in the form of instructions. More recent proprietary chat interface services such as ChatGPT (OpenAI, 2023), Gemini (Google, 2024), and BingChat are accelerating the exploration of human curiosity.
@@
LLMs’ long responses consist of innumerable combinations of relevant topics, as there is no gold standard in long-form text (Krishna et al., 2021; Xu et al., 2022, 2023). However, LLMs often fail to provide tailored responses when users’ queries get complicated. For instance, a user can request LLMs to provide information on a topic, Generative AI, especially focusing on historical context without describing recent trends (since the user already understands the latest news), but it is too complicated for LLMs to generate meticulously-composed responses. Since humans have their own background knowledge and curiosity level, we define these queries that constrain the information coverage of certain topics as coverage-conditioned (C2) queries. Specifically, our scenario conceptualizes “coverage” by assuming that users can request LLMs to include or exclude a specific subtopic within long-form responses.
Figure 1: Unlike the existing RAG systems that simply rewrite or decompose the given query, our QPlanner, trained with QTree, first explores possible information coverage. A subset of queries is selected as an outline to retrieve relevant passages and generate long-form response. By leveraging retrieved passages, we prompt LLM to generate the final response tailored to user interests.
There have been several studies to generate responses following users’ instructions (Ouyang et al., 2022; Wang et al., 2022; Wei et al., 2022a; Taori et al., 2023; Asai et al., 2023). Among these trends, we focus on query outlining, which is a sequence of queries representing intermediate subtopics within a long text, demonstrating its effectiveness in long-form story generation (Fan et al., 2018; Sun et al., 2022; Yang et al., 2022, 2023; Wang et al., 2024; Shao et al., 2024; Lee et al., 2024). On top of that, we further raise a question: “How to generate customized outlines for long-form responses in the retrieval-augmented setup?” Well-generated outlines should align with user perspectives (Shao et al., 2024) and show improved performance in downstream tasks, such as document retrieval and response generation. In our work, outlines consisting of relevant topics to C2 queries are expected to be utilized both as better search queries and intermediate content drafts, thereby improving the end-to-end retrieval-augmented generation (RAG) performance.
To explore the way to generate customized outlines, we propose QTree, 10K hierarchical sets of information-seeking queries interpreting user queries of existing datasets with diverse perspectives. Hierarchies of QTree are comprised based on the abstraction level of the main topic. For example, as illustrated in Figure 1, hierarchical topics related to processes after pretraining (i.e., Finetuning and RLHF) are selected among various viewpoints in LLMs. We assume that the selected subset of QTree performs a role as an outline for document search and response generation.
To extract a subset of queries following C2 query from QTree, we construct QPlanner, an on-premise autoregressive 7B language model generating customized and preferable outlines without human intervention. Specifically, QPlanner generates outlines to arbitrary C2 queries by learning the outline-extracting capabilities of LLMs. We evaluate the performance of QPlanner by measuring whether generated outlines follow C2 queries, both with automatic and human evaluation. Moreover, we verify that training QPlanner with preference alignment can further improve the outline quality and end-to-end RAG output. Our contributions are summarized as follows:
  • We explore the effectiveness of query outlining in coverage-conditioned (C2) scenarios by constructing QTree, 10K hierarchical sets of queries containing various perspectives.
  • We train QPlanner, an autoregressive 7B language model that generates outlines for long-form responses customized to C2 queries.
  • By conducting automatic and human evaluation, we verify the effectiveness of alignment training for outlining C2 queries and analyze how the aligned outlines help generate long-form responses, focusing on the RAG task.
@@ 2.1 Query Modification with LLMs
Integrating retrieval systems with LLMs has become crucial in enhancing retrieval-augmented systems, where query reformulation plays a pivotal role in improving information retrieval outcomes. Recent advancements involve prompting LLMs to supply detailed information, such as expected documents or pseudo-answers, for query expansion (Wang et al., 2023; Jagerman et al., 2023). Additionally, there’s a growing interest in utilizing reward signals to fine-tune the query reformulation model, further optimizing search results based on the ranking of retrieved documents (Ma et al., 2023; Yoon et al., 2024). Recent efforts have also focused on breaking down complex questions into simpler subqueries, leveraging the decomposition for more accurate retrieval and response generation (Khot et al., 2023). Additionally, a line of studies (Yao et al., 2023; Press et al., 2023; Khattab et al., 2023) merge the concept of Chain-of-Thoughts (CoT) (Wei et al., 2022b) with interaction through retrieval systems. Furthermore, most recent studies (Oh et al., 2024; Weller et al., 2024) introduce the inclusion of instructional scenarios in the information-seeking process.
Figure 2: The overview of our framework. [Step 1] Base query ($q_{base}$) is decomposed into queries with diverse viewpoints (QTree), preceded by generating coverage query ($q_{cov}$). [Step 2] After C2 candidate outlines are extracted, an LLM evaluates each outline and selects the best-scored one. [Step 3] Utilizing this dataset, QPlanner is trained to sequentially generate its own QTree and preferred outline by taking the C2 query as an input.
While having similar query modification strategies (e.g., query decomposition or instructional scenarios), we consider query outlining as an effective strategy before generating long-form responses. Previously, a hierarchical structure has shown remarkable performance on story generation task (Fan et al., 2018). Yang et al. (2022, 2023) have also emphasized the importance of outline for structured story generation. More recently, Shao et al. (2024) have explored the role of the outline as a medium to present diverse perspectives and constructed outlines by iterative conversational processes. Furthermore, Lee et al. (2024) have improved the quality of free-form writing by outline augmentation. Above these studies, we position our work as the automatic exploration of a better outline for retrieval-augmented long-form generation in C2 scenarios.
@@ 2.2 Evaluation of Long-form Responses
Evaluating long-form responses of LLMs presents a significant challenge due to the subjective and complex nature of such assessments. Recent research (Krishna et al., 2021; Xu et al., 2022, 2023) highlight the limitations of current automated metrics in accurately assessing long-form texts, underscoring the necessity for detailed, fine-grained evaluation methods. Some holistic approaches consider a wide array of model capabilities, introducing multi-metric evaluation settings (Liang et al., 2023; Gehrmann et al., 2023; Shevlane et al., 2023; Ye et al., 2024). Additionally, the push towards fine-grained atomic evaluation settings for tasks like fact verification and summarization indicates a growing interest in detailed assessments (Min et al., 2023; Krishna et al., 2023). Recent research has also explored model-based evaluation, where learned models determine automated scores (Yuan et al., 2021; Liu et al., 2023; Kim et al., 2024). While these studies consider long-form responses themselves as the target to evaluate, we evaluate both outlines and responses by following insights from Kellogg (1988) where outlines alleviate the cognitive overload of reading long text (i.e., serving as an effective content draft) while representing the core structure of long-form writing.
@@ 3 Framework
3.1 Definition
Throughout our work, we call QTree as a hierarchical set of queries derived from a single user query. Formally, QTree consists of a hierarchical query sequence $\left\{q_{1},q_{2},...,q_{11},...,q_{ijk}\right\}$, where the length of each query annotation denotes the hierarchical level within QTree. For example, $q_{213}$ refers to the query located at the third position at depth-3, following the path from the second position at depth-1 to the first position at depth-2. We set the depth and the width of QTree as three, respectively (i.e., 3+9+27=39 queries exist in each QTree). Additionally, as exemplified in Figure 1, we define C2 query as the concatenation of user’s original query (base query; $q_{base}$) and additional query constraining information coverage (coverage query; $q_{cov}$), formalized as $C^{2}=[q_{base};q_{cov}]$.
@@ 3.1 Definition
ExpertQA (Malaviya et al., 2023) Expert 1,741 100
@@ 3.2 Overview
Figure 2 illustrates the overall pipeline of our framework. Followed by $q_{base}$ collection (Section 3.3.1), we construct QTree (Section 3.3.2) and generate $q_{cov}$ (Section 3.3.3). Since we intend to generate C2 queries dealing with users’ various requests, QTree has to interpret $q_{base}$ with diverse perspectives. For example, a $q_{base}$What is Generative AI?" can be answered through various viewpoints, including its latest trend, historical context, and application across different fields. By utilizing generated $q_{cov}$, candidate outlines for each C2 query can be parsed (Section 3.4.1), preceded by the evaluation of each outline for selecting the best one (Section 3.4.2). The following subsections detail the construction procedure, and all used prompts are provided in Appendix E.
@@ 3.3.1 Base Query ($q_{base}$) Collection
We first collect $q_{base}$ that requires long-form content composition to respond. Specifically, we utilize two Wikipedia-based long-form question answering datasets – ASQA (Stelmakh et al., 2022) and Longform (Köksal et al., 2024), and one from expert domain – ExpertQA (Malaviya et al., 2023), as demonstrated in Table 1. For the test set, we sample 100 test queries for each dataset. By leveraging LLMs [^1] We fix the LLM with the gpt-4-0125-preview of OpenAI (OpenAI, 2023), with a temperature of 1.0, throughout this work., we construct C2 queries by combining these 10K $q_{base}$ with corresponding $q_{cov}$ (see Section 3.3.3). Since there are a few noisy queries, we slightly modify and filter existing queries. We describe the detailed preprocessing in Appendix A.
@@ 3.3.2 QTree Construction
Prior to generating $q_{cov}$, we decompose $q_{base}$ into diverse queries as a tree structure (i.e., QTree). The purpose of constructing QTree for each $q_{base}$ is to unfold the scope of information within parametric knowledge of LLMs. This structured graph also enables effective instruction generation (will be detailed in the next subsection) according to the hierarchy of abstractiveness. We assume queries in deeper depth present more concrete subtopics. In other words, queries in leaf nodes contain the most detailed queries of a specific subject. Table 6 in Appendix C illustrates an example of QTree.
@@
Checking the quality of QTree is relatively easier than unformatted text sequence. In the query decomposition stage, we ensure that QTree contains a predefined number of nodes in each depth and width and does not overlap each other. This can be simply done by heuristically inspecting the structured output. We also prompt the LLM to generate queries as question-format [^2] We intend every query contains question mark, but we do not explicitly filter queries with non-question format as questions can be expressed in various forms..
@@ 3.3.3 Coverage Query ($q_{cov}$) Generation
To remind, $q_{cov}$ constrain coverage within the wide range of information. Therefore, generating $q_{cov}$ from QTree requires selecting specific viewpoint to cover. However, solely relying on LLMs’ parametric knowledge does not guarantee the diversity of realistic situations. We therefore adopt the following two concepts to assist $q_{cov}$ generation.
  • Background Query: Understandably, users’ interests attribute to their background knowledge. With this in consideration, we randomly select a single query from QTree, which will be a starting point for simulating users’ curiosity. We define this query as background query. The specificity of the background query also differs according to the depth of the selected query. In other words, it can be inferred that users who are seeking information of deeper queries possess a more detailed recognition of the relevant subjects.
  • Intent Operation: While users consider a particular subject to ask, they may choose whether the content should be addressed within the responses. We conceptualize user intent through a binary operation (i.e., Inclusion, Exclusion), thereby facilitating the generation of $q_{cov}$ that explicitly request the inclusion/exclusion of the subtopic on the background query.
In practice, we prompt LLM to generate $q_{cov}$ by combining randomly selected background query from QTree with intent operation. As shown in Table 2, combinations of background query and intent operation yield diverse $q_{cov}$ for each $q_{base}$. Especially, in Exclusion operation, we find that relatively harder requirements (e.g., avoiding one topic but focusing on another topic) are given. Therefore, it can be assumed that $q_{cov}$ with Exclusion operation requires harder scenarios (we analyze this in Appendix D). In addition, we generate five $q_{cov}$ per each $q_{base}$ and choose one if corresponding outlines are correctly parsed (will be further described in Section 3.4.1).

### v2

@@ Abstract
Interactions with large language models (LLMs) often yield long and detailed responses, leveraging both parametric knowledge and retrieval-augmented generation (RAG). While these responses can provide rich insights, they often include redundant or less engaging content not aligned with user interests. This issue becomes apparent when users specify particular subtopics to include or exclude – termed coverage-conditioned (C2) queries – as LLMs often struggle to provide tailored responses. To address this challenge, we investigate the role of query outlines, sequences of subqueries designed to guide LLMs in generating responses that meet specific user requirements. To systematically create and evaluate these outlines, we introduce QTree, a dataset of 10K hierarchical sets of information-seeking subqueries that define structured boundaries for outline creation and evaluation in C2 scenarios [^1] Our resources are available at https://github.com/youngerous/qtree.. Additionally, we develop QPlanner, a 7B language model trained to generate customized outlines within boundaries of QTree. We evaluate the effectiveness of the generated outlines through automatic and human judgements, focusing on their impact within retrieval-augmented generation (RAG) systems. Experimental results demonstrate that QPlanner, especially when trained with alignment techniques like DPO, generates higher-quality outlines that better fulfill diverse user needs.
Learning to Explore and Select for Coverage-Conditioned Retrieval-Augmented Generation
@@ 1 Introduction
Figure 1: QTree constrains the range of available outlines for the user’s C2 query, and tailored outlines satisfying the requirement of C2 query are selected for RAG downstream tasks.
Recent advancements of large language models (LLMs) have enabled them to provide long and detailed responses by leveraging their parametric knowledge. As these models improve, human-machine interaction interfaces (e.g., chat and acoustic interfaces) – which have been studied for a long time (Levin et al., 2000) – have become more sophisticated, allowing users to request highly specific and personalized information. Proprietary chat services such as ChatGPT (OpenAI, 2023), Gemini (Google, 2024), and BingChat have further accelerated the exploration of personalized information. Additionally, retrieval-augmented generation (RAG) methods are being adopted to enhance the relevance and timeliness of LLM responses by integrating external knowledge.
@@
Despite these advancements, LLMs often struggle with delivering tailored responses when faced with complex user queries. For instance, a user might request LLMs to provide information on Generative AI, focusing specifically on its historical context while excluding recent trends. Crafting such meticulously composed responses is difficult for LLMs for two reasons: (1) LLMs’ long-form outputs can contain innumerable combinations of relevant topics, and (2) there is no established gold standard for long-form text generation (Krishna et al., 2021; Xu et al., 2022, 2023). Recognizing this, we first define queries that constrain the information coverage of certain topics as coverage-conditioned (C2) queries, where “coverage” refers to the user’s intent to instruct LLMs to include or exclude specific subtopics within their responses. These C2 queries especially pose challenges in constructing long-form RAG responses as they require selective document retrieval as well.
To improve LLM responses for users’ complex queries, there have been works on query outlining, creating sequences of intermediate subtopics to guide long-form responses. Query outlining has been effective in areas like long story generation (Fan et al., 2018; Sun et al., 2022; Yang et al., 2022, 2023; Wang et al., 2024; Shao et al., 2024). However, generating high-quality outlines that address complex queries like C2 queries remains challenging, as there is no systematic approach for creating and evaluating such outlines.
With the concepts of C2 query and query outlining in place, we pose two key research questions:
  1. RQ1.
    How can we create and evaluate better outlines for C2 queries?
  2. RQ2.
    Can these outlines improve RAG systems by serving as search queries and content drafts?
To address RQ1, we present QTree, a dataset comprising 10K hierarchical sets of information-seeking subqueries (with 39 subqueries in each set) that interpret user queries with diverse perspectives, facilitating the exploration and selection of appropriate outlines for C2 queries. The hierarchies in QTree are organized according to the abstraction level of the main topic, defining tangible boundaries of available outlines. For example, as illustrated in Figure 1, hierarchical subtopics related to processes after pretraining (i.e., Fine-tuning and RLHF) are selected as proper outlines for RAG response among various viewpoints on the topic of Training LLMs, following the requirements of the C2 query. In contrast, less relevant subtopics in QTree (e.g., Pretraining LLMs) will not be a desirable outline for the C2 query. By leveraging QTree, we can systematically create and judge outlines for long-form responses, ensuring that they align with the user’s coverage constraints.
Regarding RQ2, we introduce QPlanner, an autoregressive 7B language model designed to generate tailored outlines within QTree’s hierarchical boundaries. We hypothesize that high-quality outlines aligned with C2 queries can improve both document retrieval and response generation in RAG systems. We also evaluate QPlanner ’s performance through both automatic metrics and human judgments, assessing the quality of the generated outlines and their impact on downstream tasks. Experimental results on C2 queries from diverse domains (i.e., Wikipedia and expert domains) demonstrate that training QPlanner with preference alignment further improves both outline quality and overall RAG performance.
Our contributions are summarized as follows:
  1. 1.
    We present QTree, a novel dataset of 10K hierarchical subquery sets that define boundaries for available outlines, facilitating the creation and evaluation of better outlines for coverage-conditioned (C2) queries (addressing RQ1).
  2. 2.
    We introduce QPlanner, an autoregressive language model designed to generate customized outlines that improve document retrieval and content generation in RAG systems (addressing RQ2).
  3. 3.
    We conduct comprehensive evaluations, including automatic metrics and human judgments, to validate the effectiveness of our approach in enhancing outline quality and RAG performance.
@@ 2 Related Work
Figure 2: The overview of our framework. [Step 1] Base query ($q_{base}$) is decomposed into subqueries with diverse viewpoints (QTree), preceded by generating coverage query ($q_{cov}$). [Step 2] After C2 candidate outlines are extracted, a judge LLM evaluates each outline and selects the best-scored one. [Step 3] Utilizing this dataset, QPlanner is trained to sequentially generate its own QTree and preferred outline by taking the C2 query as an input.
@@ 2.1 Query Modification with LLMs
Integrating retrieval systems with LLMs has become crucial, with query modification playing a pivotal role in improving information retrieval outcomes. Recent advancements focus on prompting LLMs to provide detailed information, such as expected documents or pseudo-answers, for query expansion (Wang et al., 2023; Jagerman et al., 2023). Furthermore, reward signals are being used to fine-tune query modification models, optimizing search results based on the ranking of retrieved documents (Ma et al., 2023; Yoon et al., 2024). Additionally, complex questions are being decomposed into simpler subqueries to enhance retrieval accuracy and response generation (Khot et al., 2023).
Among various query modification strategies, query outlining stands out as an effective method for generating long responses. Outlining techniques have been primarily studied in tasks such as story generation (Fan et al., 2018). Yang et al. (2022, 2023) have also emphasized the importance of outline for narrative generation, while Shao et al. (2024) explored the outline as tools for presenting diverse perspectives through iterative conversational processes. More recently, Lee et al. (2024) improved free-form writing with outline augmentation. However, despite these advances, it has seen less attention in retrieval-augmented contexts. In addition, none of these studies systematically evaluate the generated outlines in complex scenarios (e.g., C2 scenarios). Our work aims to address this gap by proposing a controlled evaluation testbed for outlines and their impact on long-form responses.
@@ 2.2 Evaluation of Long-form Responses
Evaluating long-form responses from LLMs presents a significant challenge due to the subjective and multifaceted nature of the task. Previous studies (Krishna et al., 2021; Xu et al., 2022, 2023) highlight the limitations of automated metrics in accurately assessing long-form texts, underscoring the need for more nuanced evaluation methods. Several approaches have emerged to tackle this issue by incorporating multi-metric evaluation frameworks (Liang et al., 2023; Gehrmann et al., 2023; Shevlane et al., 2023; Ye et al., 2024), as well as task-specific metrics for fact verification and summarization (Min et al., 2023; Krishna et al., 2023). Recent research has also investigated model-based evaluations where learned models are used to generate automated scores (Yuan et al., 2021; Liu et al., 2023; Kim et al., 2024a).
While most of these studies focus solely on the evaluation of long-form responses, we extend this by evaluating both the outlines and responses they guide. Drawing from insights in cognitive psychology (Kellogg, 1988), we argue that outlines alleviate the cognitive overload for readers, functioning as effective content drafts and providing the core structure for long-form writing.
@@ 3 Framework
3.1 Background
We refer to QTree as a tree-shaped hierarchical set of subqueries (defining “subquery” as each node in QTree) derived from a single user query. We set both the depth and the width of QTree at three levels (i.e., 3+9+27=39 subqueries in each QTree). Additionally, as illustrated in Figure 1, we define C2 query as the concatenation of the user’s original query (base query; $q_{base}$) and additional coverage-constraining query (coverage query; $q_{cov}$), represented as $C^{2}=[q_{base};q_{cov}]$.
@@ 3.1 Background
ExpertQA (Malaviya et al., 2024) Expert 1,741 100
@@ 3.2 Overview
Figure 2 illustrates the procedural framework, including the construction of QTree and QPlanner. Followed by collecting $q_{base}$ (Section 3.3.1), we construct QTree (Section 3.3.2) and generate $q_{cov}$ (Section 3.3.3). Generated C2 queries (i.e., $q_{base}$ and $q_{cov}$) are then utilized to select candidate outlines. For example, the answer to a $q_{base}$What is Generative AI?" can contain diverse perspectives, including its latest trend, historical context, and application across different fields. Within available outlines that guide to satisfying answers, our goal is to obtain the outline that follows $q_{cov}$Tell me about its historical context". Therefore, within the range of QTree, we parse candidate outlines for each C2 query (Section 3.4.1), preceded by the evaluation for selecting the optimal outline (Section 3.4.2). The following subsections detail the procedural generation, and all used prompts are provided in Appendix E.
@@ 3.3.1 Base Query ($q_{base}$) Collection
We first collect $q_{base}$ that requires long-form content composition to respond. Specifically, we employ two Wikipedia-based long-form question answering datasets – ASQA (Stelmakh et al., 2022) and Longform (Köksal et al., 2024), and one from expert domains – ExpertQA (Malaviya et al., 2024), as demonstrated in Table 1. For the test set, we sample 100 test queries for each dataset. By leveraging LLMs [^2] We use gpt-4-0125-preview of OpenAI (OpenAI, 2023) with a temperature of 1.0, throughout this work., we construct C2 queries by combining these 10K $q_{base}$ with corresponding $q_{cov}$ (see Section 3.3.3). We slightly modify and filter a few $q_{base}$ containing noises, described in Appendix A.
@@ 3.3.2 QTree Construction
Prior to generating $q_{cov}$, we decompose $q_{base}$ into diverse subqueries as a tree structure (i.e., QTree). The purpose of constructing QTree for each $q_{base}$ is to unfold the scope of information within parametric knowledge of LLMs. This structured graph also enables effective instruction generation (will be detailed in the next subsection) according to the hierarchy of abstractiveness. Subqueries in deeper depth present more specific subtopics. Table 6 in Appendix C illustrates an example of QTree.
@@
In the query decomposition stage, we ensure that QTree contains a predefined number of subqueries (i.e., three) in each depth and width and does not overlap each other. This can be simply done by heuristically inspecting and comparing the structured output.
@@ 3.3.3 Coverage Query ($q_{cov}$) Generation
To remind, the role of $q_{cov}$ is to specify certain subtopics to address (i.e., include or exclude) within a broad range of information. Therefore, generating $q_{cov}$ from QTree requires selecting a specific viewpoint to cover. However, solely relying on LLMs’ parametric knowledge does not guarantee the diversity of realistic situations. We therefore adopt the following two concepts to assist in generating $q_{cov}$.
  • Background Subquery: Understandably, asking for specific knowledge means that users are recognizing the knowledge itself. With this in consideration, we randomly select a single subquery from QTree, which will be the knowledge users are aware of. We define this subquery as background subquery. The specificity of the background subquery differs according to the depth of the selected query.
  • Intent Operation: While considering a particular subject to ask, users may choose whether the content should be addressed within the responses. We conceptualize user intent through a binary operation (i.e., Inclusion, Exclusion), thereby facilitating the generation of $q_{cov}$ that explicitly request the inclusion/exclusion of the subtopic on the background subquery.
In practice, we prompt LLM to generate $q_{cov}$ by combining a randomly selected background subquery from QTree with intent operation [^3] Although we use background subquery to generate $q_{cov}$ in this section, it is also used to construct baselines. Refer to Section 4.2.. As demonstrated in Table 2, combinations of background subquery and intent operation yield diverse $q_{cov}$ for each $q_{base}$. Especially, requirements of $q_{cov}$ with Exclusion operation are more complicated (e.g., avoiding one topic but focusing on another topic) than Inclusion. We analyze the performance difference according to intent operations in Appendix D. We sample five preliminary $q_{cov}$ per each $q_{base}$ and finally choose one if corresponding three candidate outlines are parsed correctly (which will be further described in Section 3.4.1).